a uniformly charged and infinitel long line having a linear charge density x is placed at distance y from pt o consider an imaginary sphere of radius r with o as  center and r>y electric flux is ?
 

Grade:12

1 Answers

Eshan
5 years ago
let the length of the rod inside the sphere bel.

From geometry,y^2+(\dfrac{l}{2})^2=r^2
Hencel=2\sqrt{r^2-y2}
Therefore the enclosed charge inside the sphere=xl=2x\sqrt{r^2-y2}
Hence the flux passing through the sphere==\dfrac{2x\sqrt{r^2-y2}}{\epsilon_0}